Find three consecutive positive odd integers such that 4 times the sum of all three equals 13 more than the product of the smaller two.
Answer by stanbon(75887)

Find three consecutive positive odd integers such that 4 times the sum of all three equals 13 more than the product of the smaller two.
--------------------------------------------------------------
1st: 2x-1
2nd: 2x+1
3rd: 2x+3
================
Eequation:
4(6x+3) = (2x-1)(2x+1) + 13
24x+12 = 4x^2-1 + 13
4x^2 -24x = 0
4x(x-6) = 0
x = 0 or x = 6
---
Positive solutions:
1st 2x-1 = 2*6-1 = 11
2nd : 13
4rd : 15
=======================
Cheers,
Stan H.